ok i am in Algebra A and i am doing weighted averages. the problem is: An alloy of metals is 25% copper.Another alloy is 50% copper. how much of each alloy should be used to make 1000 grams of an alloy that is 45% copper?
Found 2 solutions by stanbon, ankor@dixie-net.com:
Answer by stanbon(48524)

An alloy of metals is 25% copper.
Another alloy is 50% copper.
how much of each alloy should be used to make 1000 grams of an alloy that is 45% copper?
---------------
Equation:
copper + copper = copper
0.25x + 0.50(1000-x) = 0.45(1000)
0.25x + 500-0.50x = 450
-0.25x = -50
x = 200 grams (amt. of 25% alloy in the mix)
1000-x = 800 grams (amt of 50% alloy in the mix)
================
Cheers,
Stan H

Answer by ankor@dixie-net.com(12692)

An alloy of metals is 25% copper.Another alloy is 50% copper. how much of each alloy should be used to make 1000 grams of an alloy that is 45% copper?
:
Let x = amt of 50% copper
It says the resulting amt is 1000 gm, therefore:
(1000-x) = amt of 25% copper
:
.25(1000-x) + .50x = .45(1000)
:
250 -.25x + .50x= 450
+.25x = 450 - 250
x = 200/.25
x = 800 gm of 50% alloy is required
then
1000 - 800 = 200 gm of 25% alloy required
:
:
Check solutions
.25(200) + .5(800) = .45(1000)
50 + 400 = 450